Problem 7-16 (Algo)

Burrito King (a new fast-food franchise opening up nationwide) has successfully automated burrito production for its drive-up fast-food establishments. The Burro-Master 9000 requires a constant 50 seconds to produce a batch of burritos. It has been estimated that customers will arrive at the drive-up window according to a Poisson distribution at an average of one every 65 seconds. To help determine the amount of space needed for the line at the drive-up window, answer the following questions.

a. What is the average waiting line length (in cars)? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

b. What is the average number of cars in the system (both in line and at the window)? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

c. What is the expected average time in the system, in minutes? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

Answer: - According to given data

Arrival rate (lambda) = 65 seconds in minutes = 60/65 = 0.9230 per minute

Service rate (mue) = 50 seconds in minutes = 60/50 = 1.2000 per minute

a) Formula to calculate average line length = lambda² / 2mue(mue - lambda)

Line length = 0.9230² / 2*1.20(1.20-0.9230)

= 0.852 / 2.4(0.2769) = 1.28 cars

b) Formula to calculate number of cars in the system = Lq + lambda/mue

We already calculated Lq = 1.28 cars

= 1.28 + 0.9230/1.200 = 2.05

Average number of cars in the system = 2.05 cars

c) Formula to calculate Average time in the system = Ls / lambda

= 2.05 / 0.9230 = 2.22 minutes